### v2

# Winemaking: Extracting Essential Insights for Efficient Threat Detection in Audit Logs

Abstract.
Advanced Persistent Threats (APTs) are continuously evolving, leveraging their stealthiness and persistence to put increasing pressure on current provenance-based Intrusion Detection Systems (IDS). This evolution exposes several critical issues: (1) The dense interaction between malicious and benign nodes within provenance graphs introduces neighbor noise, hindering effective detection; (2) The complex prediction mechanisms of existing APTs detection models lead to the insufficient utilization of prior knowledge embedded in the data; (3) The high computational cost makes detection impractical.
To address these challenges, we propose Winemaking, a lightweight threat detection system built on a knowledge distillation framework, capable of node-level detection within audit log provenance graphs. Specifically, Winemaking applies graph Laplacian regularization to reduce neighbor noise, obtaining smoothed and denoised graph signals. Subsequently, Winemaking employs a teacher model based on GNNs to extract knowledge, which is then distilled into a lightweight student model. The student model is designed as a trainable combination of a feature transformation module and a personalized PageRank random walk label propagation module, with the former capturing feature knowledge and the latter learning label and structural knowledge. After distillation, the student model benefits from the knowledge of the teacher model to perform precise threat detection. Finally, Winemaking reconstructs attack paths from anomalous nodes, providing insight into the attackers’ strategies. We evaluate Winemaking through extensive experiments on three public datasets and compare its performance against several state-of-the-art IDS solutions. The results demonstrate that Winemaking achieves outstanding detection accuracy across all scenarios and the detection time is 1.4 to 5.2 times faster than the current state-of-the-art methods.
Threat Detection, Host Provenance, Knowledge Distillation
copyright: noneccs: Security and privacy Intrusion detection systemsjournalyear: 2023copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2023; April 30-May 4, 2023; Austin, TX, USAbooktitle: Proceedings of the ACM Web Conference 2023 (WWW’23), April 30-May 4, 2023, Austin, TX, USAprice: 15.00doi: 10.1145/3543507.3583336isbn: 978-1-4503-9416-1/23/04
1. Introduction
Advanced Persistent Threats (APTs)(Alshamrani et al., 2019) represent a complex form of cyber attack characterized by high stealth and strong targeting. In these attacks, perpetrators gain unauthorized access to a victim’s machine through methods such as network or software backdoors, and persist for extended periods to steal sensitive data or take control of the target machine. APTs have already infiltrated many highly secured large enterprises and institutions that are based on web services, causing substantial financial losses(APT, 2020; Khaleefa and Abdulah, 2022; Sharma et al., 2023). Notable examples include the Equifax(Equ, 2017) breach, which resulted in a record number of user data being stolen, and the SolarWinds(Sol, 2020) attack, which had a vast scope and severe impact.
To combat these sophisticated APTs attacks, host-based Intrusion Detection intrusion detection systems (IDS) have become a widely deployed defense mechanism. However, with the evolving nature of attack techniques and the growing scale of threats, traditional IDS solutions can no longer meet the demands of the changing threat landscape. Currently, provenance -based(Li et al., 2021) detection methods are considered effective for capturing APTs. These methods apply system audit logs and structure system entities (such as processes, files, and network flows) into a graph structure known as a provenance graph. Based on how audit logs are utilized, these detectors can be divided into three categories: statistics-based detection(Hassan et al., 2019, 2020b; Liu et al., 2018) quantifies the suspiciousness of audit logs by analyzing the rarity in the provenance graph; rule-based detection(Hassan et al., 2020a; Hossain et al., 2017; Milajerdi et al., 2019) matches audit logs with attack patterns using expert security knowledge bases; and learning-based detection(Jia et al., 2024; Zengy et al., 2022; Cheng et al., 2023; Wang et al., 2022; Han et al., 2020; Rehman et al., 2024) employs machine learning techniques to learn from the provenance graph, identifying abnormal system behaviors and attack patterns. Among these approaches, learning-based detection has been considered the most promising in recent years. However, we have observed several persistent challenges that impact real-world detection:
  • Neighbor Noise: In provenance graphs constructed from audit logs, malicious nodes often exhibit long-distance, multi-hop distributions, meaning that a complete attack can involve more benign system entities. The interaction between malicious and benign nodes generates interference and confounds the classifier in detection systems. For node-level classification tasks, the neighboring nodes adjacent to the target detection node introduce dense neighbor noise, which can obscure true anomalies, leading to false positives or missed detections, and ultimately reducing detection accuracy.
  • High Computational Cost: Graph-based algorithms enable learning from provenance graphs to capture the complex relationships within them. While these methods can achieve impressive detection performance, they come at the cost of significant memory and time overhead. As a result, these approaches lose the capability for real-time detection and can only function as offline systems to analyze graph data(Zengy et al., 2022; Wang et al., 2022).
  • Insufficient Utilization of Prior Knowledge: Recent APTs detection methods(Jia et al., 2024; Zengy et al., 2022; Cheng et al., 2023) primarily rely on large-scale graph algorithms, such as GNNs, to extract structural and node feature information from graphs. However, the prediction mechanism of GNNs is highly complex, as it tightly integrates graph topology, node features, and projection matrices. This entanglement makes it challenging to clearly interpret the relationships between these factors within the model. This complexity prevents the effective utilization of prior knowledge in terms of labels, features, and structure(Li et al., 2019; Wang and Leskovec, 2020).
Thus, we propose Winemaking, a lightweight APTs detection method that achieves both high detection accuracy and computational efficiency. Specifically, Winemaking incorporates the following key functionalities: (1) Winemaking constructs a provenance graph from audit logs that include various APTs attacks and uses Word2Vec to assign initial features to nodes. To address the issue of neighbor noise, Winemaking designs a graph Laplacian regularization-based graph signal denoising method tailored to the provenance graph. This algorithm smooths node signals without disrupting the original graph topology. (2) Winemaking introduces a knowledge distillation framework where, during training, knowledge from a pre-trained large GNN teacher model is distilled into a lightweight student model. The lightweight student model is then used for anomalous node detection, significantly reducing detection costs. (3) For the student model, Winemaking proposes a new hybrid mechanism that combines the Feature Transformation (FT) mechanism with a personalized PageRank random walk label propagation (PRL). This allows more effective learning of prior knowledge from features, labels, and structures while utilizing the GNN knowledge from the teacher model. (4) After detecting anomalous nodes, Winemaking uses a community division algorithm to identify malicious communities.
We performed a comprehensive evaluation of Winemaking using datasets widely adopted by the research community, including StreamSpot(The, 2016), Unicorn Wget(Han et al., 2020), and DARPA-E3(DAR, 2020). We first analyzed its detection performance and compared it against various existing baselines, demonstrating that Winemaking consistently outperforms almost all these systems under the same evaluation metrics. Furthermore, we analyzed Winemaking’s detection costs and explored its potential as a real-time detection system. In addition, hyperparameter analysis and ablation experiments on Winemaking highlight the irreplaceability of its key parameters and main components. Lastly, we verified Winemaking’s inherent robustness against adversarial attacks. The primary contributions of our work are as follows:
Figure 1. Attack scenario from DARPA E3 CADETS. Green indicates benign entities and red signifies malicious entities. R = Read, W = Write, O = Open, E = Execute, S = Send and Rc = Receive.
[…]
A figure illustrating four scenes.
  • We propose Winemaking, a general APTs detection method based on knowledge distillation, capable of transferring GNN knowledge from a large teacher model to a lightweight student model, thereby reducing resource and time costs during detection.
  • We address the issue of neighbor noise in provenance graphs by designing a graph Laplacian regularization-based graph signal denoising method that smooths and denoises signals tailored for provenance graphs.
  • We examine the problem of insufficient utilization of prior knowledge caused by complex prediction mechanisms in current approaches and design a lightweight student model to effectively learn and leverage prior knowledge.
  • We comprehensively evaluate Winemaking ’s performance by implementing it on three widely used datasets and conducting various experiments to verify its effectiveness. Experimental results demonstrate that Winemaking outperforms most existing methods while incurring lower detection costs.
2. Attack scenarios and Threat model
2.1. Motivating Example
This is an attack case from the DARPA E3 CADETS: where the attacker establishes a connection with the Nginx server via IP address 81.49.200.166 and gains shell access. Using the shell, the attacker downloads a malicious payload to /tmp/vUgefal and executes it. The process vUgefal then moves laterally to 61.167.39.128, completes the infection, and further injects sshd into the system, downloading another payload to /var/log/devc. Figure 1 illustrates our motivating example, where three types of nodes represent system entities, and the lines indicate the directional interactions between them.
Table 1. Limitations of existing provenance-based intrusion detection systems.
Model
Neighbor
Denoising
Lightweight
Models
Utilization of
Prior Knowledge
Attack
Reconstruction
Granularity
Winemaking (ours) Node
MAGIC(Jia et al., 2024) Node
KAIROS(Cheng et al., 2023) Node
ThreaTrace(Wang et al., 2022) Node
ShadeWatcher(Zengy et al., 2022) Edge
Unicorn(Han et al., 2020) Graph
2.2. Limitations to Existing Solutions
Learning-based detection methods are currently the most promising in the field of threat detection. Their core principle is to build anomaly models by learning from large datasets and then infer whether anomalous behavior represents a malicious attack. Some of the most notable recent works include: MAGIC(Jia et al., 2024), which performs efficient detection by masking the graph and reducing costs; KAIROS(Cheng et al., 2023), which identifies four dimensions of intrusion detection systems and uses an encoder-decoder architecture to detect attacks and reconstruct attack chains; THREATRACE(Wang et al., 2022), the first detector to propose behavior pattern construction for each type of node in the provenance graph to achieve node-level detection; ShadeWatcherShadeWatcher(Zengy et al., 2022), which introduces recommendation system techniques into provenance graph detection and achieves edge-level detection; Unicorn(Han et al., 2020), which builds a provenance graph enriched with semantics and historical information to capture long-term stealthy attacks. Although these works demonstrate excellent performance in threat detection, they still face limitations when tested in real-world environments. These limitations are what Winemaking aims to address, as shown in Table 1.
  • Neighbor Denoising: Provenance graphs constructed from audit logs contain rich information, with nodes and edges representing different meanings. However, current approaches seem too eager to move directly to graph representation learning without adequately addressing the noise in the initial graph. We observed that only Magic acknowledged this issue.
  • Lightweight Models: One of the major challenge with current log provenance algorithms lies in the large size and difficulty of deployment of the models. Most existing approaches rely on GNN models, which, due to their multi-layer architectures and large parameter inputs, result in bloated models. Furthermore, the high memory and time overhead makes these detectors impractical and limits their deployment in real-world environments. Notably, only Magic(Jia et al., 2024) has optimized performance costs to address this issue.
  • Utilization of Prior Knowledge: Labels, node features, and graph structures offer rich prior knowledge that can significantly enhance model performance. Label information helps the model capture the true class relationships of system entities. Node features reflect the attribute differences among system entities, and the graph structure reveals the topological relationships and latent patterns between interactions. However, existing APTs detection methods often over-rely on complex network topologies and automated feature learning, overlooking the explicit prior knowledge embedded in these components.
  • Attack Reconstruction: A simple and complete attack chain helps security analysts better understand the attack behavior. However, from a large number of anomalous nodes detected at the node level, we need to reconstruct these attack paths based on the dependency relationships between kernel objects, as exemplified by KAIROS, to enable analysts to respond quickly.
  • Granularity: Different detection granularities yield different levels of attack discovery. For example, graph-level granularity, as in Unicorn, is easy to partition but lacks the precision to pinpoint specific malicious entities in the graph. Edge-level granularity, as in ShadeWatcher, offers more detailed insights into specific edge relationships but incurs higher detection costs. Node-level granularity methods like Magic, KAIROS, and THREATRACE, can target specific entities without requiring detailed learning of edge relationships.
2.3. Threat Model
To build a more comprehensive threat model, we make several assumptions about attacker behavior and system characteristics. We consider APTs attackers, who aim to infiltrate a target host through network channels, pre-installed backdoors, or software vulnerabilities. This process is often highly stealthy and can persist for an extended period. During the intrusion, attackers attempt to blend their malicious activities with legitimate background data to obscure their intent. For instance, they may use Living-off-the-land techniques(Ongun et al., 2021), injecting malicious code into legitimate processes, which then spawn new legitimate processes to continue their malicious actions.
Although the attackers’ behavior is highly covert, audit logs capture complete, traceable evidence and footprints of their activities. As in most of the existing works in this field(Wang et al., 2022; Han et al., 2020; Rehman et al., 2024), we assume that the underlying operating system, audit framework, and system hardware are intact and trustworthy(Bates et al., 2015; Pasquier et al., 2017). Additionally, we assume that attackers cannot directly tamper with the contents of the audit logs, ensuring that the provenance graph constructed by Winemaking remains reliable(Paccagnella et al., 2020).
3. Methodology
Figure 2. The framework of Winemaking.
In this section, we provide a detailed overview of Winemaking’s overall design framework and the design of each module. The framework consists of the following components: ①Graph Construction. ②Log Distillation. ③Threat Detection. ④Attack Reconstruction. The overall framework is illustrated in Figure 2.
3.1. Graph Construction
3.1.1. Provenance Graph Construction
Winemaking processes host audit logs from various sources(such as Windows ETW, Linux Audit and CamFlow(Pasquier et al., 2017)) and transforms them into a graph structure known as a Provenance Graph (PG). In this graph, nodes represent system entities, and edges denote the interactions between these entities.
Since system logs contain rich attributes related to various system entities, Winemaking encodes these attributes into a vector space to serve as initial input for model learning. We use Word2Vec(Ma and Zhang, 2015), a neural network-based technique effective at learning word vector representations, which generates dense, low-dimensional vectors for each word. In our cataloging process, we use attributes such as file paths for files, process names for process nodes, system entity types, and network IP addresses for sockets to generate feature vectors. These are combined with semantic attributes and system call types within one hop of the neighbors to form sentences. During training, Winemaking captures semantic relationships between words within a fixed length, outputting embedded vectors as the raw signal for subsequent modules.
3.1.2. Neighbor Denoising
Given that in the provenance graph constructed from audit logs, malicious nodes (i.e., system entities manipulated by attackers) are rare, and adversarial activities—such as deceptive access to benign nodes or the manipulation of compromised benign nodes to perform misleading actions—are common, there exists a significant issue of neighbor noise for anomaly detection tasks. To address this, Winemaking incorporates a graph Laplacian regularization-based signal denoising method, which is designed to leverage the graph’s topological structure to smooth signals and reduce the impact of noise.
Specifically, Winemaking defines the weights based on the similarity and connection relationships between entity nodes and combines them into a weight matrix $W$. By calculating the node degree matrix $D$, we can then derive the Laplacian matrix $L=D-W$. This enables the formulation of a convex quadratic optimization problem to minimize noise and smooth signals as follows:
(1) $\min_{\mathbf{x}}\left\{\frac{1}{2}\|\mathbf{x}-\mathbf{x}^{(0)}\|^{2}+\frac{\gamma}{2}\mathbf{x}^{\top}\mathbf{L}\mathbf{x}\right\}.$
Where |||||$\cdot$||| | ⋅ | | represents the Euclidean norm, $x$ is the raw signal, and $\gamma$ is the regularization parameter that controls the trade-off between data fidelity and signal smoothing. To achieve this, we solve the following linear system to find the optimal solution and obtain the denoised node signals:
(2) $(\mathbf{I}+\gamma\mathbf{L})\mathbf{x}=\mathbf{x}^{(0)},\quad\mathbf{x}=(\mathbf{I}+\gamma\mathbf{L})^{-1}\mathbf{x}^{(0)}.$
After the neighbor denoising process, we obtain denoised and smoothed signal vectors. It is worth noting that Winemaking’s neighbor denoising algorithm does not alter the graph’s topology in any way, ensuring that no new noise is introduced due to changes in the structure. This preserves the integrity of the original graph topology while effectively reducing the noise, allowing for more accurate and reliable detection in subsequent steps.
3.2. Log Distillation
In this section, we describe the knowledge distillation mechanism designed for Winemaking ’s log provenance graph, which explains the training process of the teacher and student models and how knowledge is distilled from the large-scale teacher model to the lightweight student model.
3.2.1. Teacher Model
The teacher model is the cornerstone of Winemaking ’s APTs threat detection system, essential for reducing detection time costs. The teacher model is pre-trained before real-time detection occurs. For Winemaking ’s teacher model, we employ various large-scale GNN models, such as GCN(Kipf and Welling, 2016), GAT(Veličković et al., 2017), GraphSAGE(Hamilton et al., 2017), and SGC(Wu et al., 2019). These GNN models are well-suited for leveraging the complex interactions between system entities found in audit logs and capturing causal connections and semantic information across entities. This characteristic is commonly utilized by the most advanced IDS today.
Unlike other IDS, Winemaking ’s teacher model is trained using supervised labels on system entities fed into the GNN model, producing both the predicted result $f_{t}hr$ and soft labels. Soft labels represent the aggregated anomaly probabilities for all test nodes and serve as an important self-supervised signal for the subsequent student model. By pre-training the teacher model on the large-scale graph topology derived from the log data and generating soft labels, Winemaking ensures robust support for real-time detection.
3.2.2. Student Model
The core requirement for Winemaking ’s student model is to compress detection time while maintaining high accuracy and effectively utilizing prior knowledge. Thus, the student model is designed to be lightweight and efficient, built on two foundational mechanisms: 1) feature transformation for individual nodes, and 2) label propagation(Iscen et al., 2019) between nodes. The former focuses on enhancing the model’s ability to fit based on each node’s intrinsic features, while the latter propagates label information across nodes.
Winemaking ’s student model integrates both mechanisms, combining Feature Transformation (FT) and Personalized PageRank Random walk Label propagation (PRL). Feature Transformation enables each node to adjust and refine its feature representation to improve detection accuracy, while PRL helps spread labels effectively across the graph to capture anomaly information in a more efficient and informed manner. This hybrid approach allows the student model to leverage the knowledge distilled from the teacher model while significantly reducing detection time and computational costs.
1. Feature Transformation We use a simple Multilayer Perceptron (MLP)(Tang et al., 2015) for feature transformation to perform nonlinear transformations on the input node features, enhancing the model’s ability to fit complex data. In general, MLP takes the raw node features as input and outputs the predicted probability for malicious nodes. Winemaking employs a two-layer MLP.
Assuming the input feature vector for a node is $x_{v}$, with an input dimension of $d$, the MLP transforms it into a feature vector $h_{v}$ as follows:
(3) $h_{v}=\sigma(W_{1}X_{v}+b_{1}).$
The prediction result from the FT module is
(4) $f_{\mathrm{FT}}(v)=\mathrm{softmax}(W_{2}h_{v}+b_{2}),$
where $W_{1}$ and $W_{2}$ are the learnable weight matrices corresponding to the first and second linear transformations, respectively. $b_{1}$ and $b_{2}$ are the bias terms, and $\sigma$ represents a non-linear activation function. The output layer passes through a softmax function to generate the final prediction.
2.Label Propagation We designed a label propagation method based on Personalized PageRank(Bahmani et al., 2010) in the form of random walks(Xia et al., 2019). This approach propagates labels from labeled nodes to unlabeled nodes by simulating random walks between them. PRL introduces a “jump-back” mechanism, which allows the label information to propagate from neighbor nodes while also having a probability of returning to the starting node, preserving more local information.
The process can be broken down into the following steps:
1) Label Initialization For any node $v\in V$, its initial label is mapped to a probability distribution based on the supervised signal from the labeled data.
2) Iterative Propagation In each iteration, PPR updates the label distribution of each node based on its current distribution and that of its neighbors. Node $v$ retains its previous label distribution from the last iteration (denoted as iteration $k$) with a probability of $1-\alpha$, meaning that it primarily relies on its own information without depending on its neighbors.
Then, with a probability $\alpha$, node $v$ receives label information from its neighboring nodes $N_{v}$. The label contribution from each neighbor $u\in{N_{v}}$ is an equally weighted average of the random walk label distribution $f_{RW}$. This ensures that each neighbor node contributes equally to the label propagation process, with all neighbors participating equally in updating the label distribution.
Finally, the updated label distribution $f_{\mathrm{RW}}^{k+1}(v)$ for node $v$ at iteration $k+1$ is the sum of its retained distribution and the propagated distribution from its neighbors:
(5) $f_{\mathrm{RW}}^{k+1}(v)=(1-\alpha)f_{\mathrm{RW}}^{k}(v)+\alpha\sum_{u\in N_{v}}\frac{f_{\mathrm{RW}}^{k}(u)}{|N_{v}|}.$
where $f_{\mathrm{RW}}^{k}(v)$ represents the label distribution of node $v$ during the $k$-th iteration, $\alpha$ is a parameter that controls the balance between jumping back and propagating the label, $N_{v}$ is the set of neighboring nodes of node $v$, and $|N_{v}|$ is the number of neighbors of node $v$, and $f_{\mathrm{RW}}^{k}(u)$ is the label distribution of node $u$.
3) Stopping Iteration The iteration stops when a fixed number of iterations is reached or when the label distribution converges. The result $f_{\mathrm{PRL}}^{k+1}(v)=f_{\mathrm{RW}}^{k+1}(v)$ is then returned.
3. Combine tasks After completing both FT and PRL, we combine them to form the complete student model task. During the learning process for each node $v$, we dynamically balance FT and PPR to make predictions. The prediction function is as follows:
(6) $f_{std}^{k+1}(v)=\beta_{v}f_{FT}(v)+(1-\beta_{v})f_{\mathrm{PRL}}^{k+1}(v)$
Where $f_{FT}(v)$ is the prediction from the Feature Transformation module, $f_{\mathrm{PRL}}^{k+1}(v)$ is the result from the label propagation (PPR), and $\beta$ is a balancing parameter that controls the contribution of each component to the final prediction.
3.2.3. Optimization Objective
The training logic of both the teacher model and the student model has been presented. In summary, the teacher model distills knowledge obtained from large-scale GNN training, while the student model is trained to mimic the soft label predictions of the pre-trained teacher model. Thus, our optimization objective is to make the student model’s prediction $f_{std}(v)$ approximate the teacher model’s soft labels $f_{thr}(v)$. The optimization goal is expressed as:
(7) $\min_{\Theta}\sum_{v\in V}distance(f_{thr}(v),f_{std;\Theta}(v))$
Assuming that the student model performs $K$ iterations, the final optimization objective $F$ can be expressed as:
(8) $\min_{\Theta}\sum_{\forall v\in V\}_{U}}\|f_{thr}(v)-f_{std;\Theta}^{K}(v)\|_{2}$
In this context, |||||$\cdot$||| | ⋅ | | represents the L2 norm, and the parameter set $\theta$ includes the balancing parameters between PPR and FT, denoted as ${\beta_{v},v\in{V}}$, the label jump-back and propagation balance parameter $\alpha$ in PPR, and the parameters $\theta$MLP from the MLP in the FT module.
Figure 3. Winemaking working modes.
[…]
A figure illustrating four scenes.
3.3. Threat Detection
Our student model operates in two modes: a training mode and a detection mode. As shown in Figure 3, during the knowledge distillation process, we transfer knowledge from the complex teacher model to the lightweight student model. In this phase, the student model is in training mode, focusing on learning the distilled knowledge from the teacher model. Once the knowledge distillation step is complete, Winemaking employs the trained student model as a detection model, which infers and classifies anomalous nodes from the provenance graph constructed from logs. At this point, the student model switches to detection mode. The detection mode workflow is as follows: we input the provenance graph (with initial features) to the student model. After lightweight $K$-layer PRL and feature transformation steps, the student model predicts an anomaly score for each node $v$. The student model efficiently passes messages and aggregates information across the input graph, classifying nodes based on the learned lightweight feature representations. The anomaly score output for each node is then compared to a pre-set threshold, and nodes exceeding this threshold are flagged as potential malicious nodes and collected into the malicious node set.
Table 2. Detection results with teacher models as GCN and GAT.
Datasets
Teacher
Student +ACC Impv.
Teacher
Student +ACC Impv.
ACC ACC PR RC F1 ACC ACC PR RC F1
StreamSpot 98.30% 99.57% 99.28% 99.81% 99.54% 1.27% 97.93% 99.82% 99.21% 99.35% 99.28% 1.89%
Unicorn Wget 98.21% 99.16% 99.63% 99.76% 99.69% 0.95% 98.27% 99.75% 99.34% 99.82% 99.58% 1.48%
DARPA CADETS 97.31% 99.77% 95.31 99.44% 97.33% 2.46% 96.56% 99.33% 95.53% 99.39% 97.42% 2.77%
DARPA TRACE 96.73% 98.35% 95.73 99.02% 97.34% 1.62% 96.22% 98.98% 95.11% 99.40% 97.21% 2.76%
DARPA THEIA 97.26% 99.94% 96.21 99.74% 97.94% 2.68% 97.43% 99.64% 96.40% 99.61% 97.98% 2.21%
3.4. Attack Reconstruction
After threat detection, security personnel may have a set of detected malicious nodes provided by Winemaking, but these nodes are often dispersed across the graph, making it difficult to directly derive the attacker’s path. To alleviate the burden on security analysts, Winemaking reconstructs a more detailed attack trace, capturing the complete attack chain without relying on prior knowledge. Specifically, Winemaking leverages Infomap(Edler et al., 2017), an algorithm based on random walks that simulates how information propagates between nodes and divides the graph into tightly connected communities. Malicious nodes and their one-hop neighbors are assigned higher weights to distinguish them from benign node connections. Next, information flow propagation is performed, where information spreads across the graph according to weighted probabilities, merging nodes to form communities.
The goal is to maximize ’intra-community propagation’ while minimizing ’inter-community propagation’. If most malicious nodes cluster within a single community, it indicates that the community may represent a core region of the APTs attack. If malicious nodes are distributed across multiple communities, it suggests that they likely serve as bridge nodes, linking different regions along the attack path.
4. Evaluation
In this section, we will conduct various experiments to validate the advantages of Winemaking and answer the following research questions:
  • RQ1: How does Winemaking ’s detection performance compare to state-of-the-art methods?
  • RQ2: How effective are the main components of Winemaking?
  • RQ3: How do Winemaking ’s hyperparameters affect its detection performance?
  • RQ4: What are the cost overheads of Winemaking as a detection system?
  • RQ5: How robust is Winemaking against adversarial attacks?
4.1. Datasets Settings
4.1.1. Datasets
StreamSpot Dataset(The, 2016): This dataset comprises provenance graphs collected by SystemTap from six controlled environments, including CNN, Download, Gmail, Vgame, YouTube, and Attack, with each environment containing 100 provenance graphs.
Unicorn Wget Dataset(Han et al., 2020): This dataset includes simulated attacks designed by UNICORN, consisting of 150 batches of log data collected by CamFlow, of which 125 batches are benign and 25 are malicious.
DARPA-E3 Dataset(DAR, 2020): This dataset is part of the DARPA TC program’s third evaluation, containing several sub-datasets. We selected the Cadets, Theia, and Trace datasets to evaluate Winemaking, using the same ground truth labels as ThreaTrace. Detailed descriptions of these datasets can be found in the appendix.
4.1.2. Metrics
We use the evaluation metrics from Unicorn(Han et al., 2020) and ThreaTrace(Wang et al., 2022) to compare performance across different granularities, enabling reasonable comparisons with other detection methods. The performance metrics include Precision (PR), Recall (RC), F1-score and Accuracy (ACC).
4.1.3. Baselines
To perform a comprehensive performance evaluation, we compared Winemaking with the best detection methods. Unfortunately, some methods were excluded due to challenges in reproducibility or incompatibility. For example, KAIROS(Cheng et al., 2023) could not be compared as it uses a time window-based approach that differs from other methods. We selected the following detection methods as our baselines:
  • StreamSpot(Jacob et al., 2008): StreamSpot detects intrusions by analyzing information flow graphs. It extracts features from the graph to learn a benign model and uses clustering methods to detect anomalous graphs.
  • Unicorn(Han et al., 2020): Unicorn utilizes graph rendering techniques to efficiently summarize long-running system executions. It classifies graphs as benign or malicious and filters anomalies.
  • Prov-Gem(Kapoor et al., 2021): Prov-Gem captures entity interactions using a unified relational-aware embedding framework, with the help of supervised signals.
  • ThreaTrace(Wang et al., 2022): ThreaTrace customizes a new GraphSAGE method to aggregate system entity nodes, enabling node-level detection.
  • Log2vec(Liu et al., 2019): Log2vec is a heterogeneous graph embedding-based method for detecting network threats within enterprises. It identifies abnormal logs using node embeddings and clustering methods.
  • DeepLog(Du et al., 2017): DeepLog leverages Long Short-Term Memory (LSTM)(Yu et al., 2019) networks to model system logs as natural language sequences and performs log-level anomaly detection.
  • MAGIC(Jia et al., 2024): MAGIC is a multi-granularity detection method that achieves efficient detection and cost reduction through graph masking techniques.
4.1.4. Implementation
Details We implemented Winemaking ’s prototype in approximately 2800 lines of Python3.11 code. For log processing, we used a log parser to convert audit log files from various sources into JSON format and then preprocess them. The development environment was PyTorch(Paszke et al., 2019), and the model implementation was supported by the Deep Graph Library (DGL)(Wang, 2019). Detailed teacher and student model parameters are provided in the A Appendix.
Table 3. Detection results with teacher models as SAGE and SGC.
Datasets
Teacher
Student +ACC Impv.
Teacher
Student +ACC Impv.
ACC ACC PR RC F1 ACC ACC PR RC F1
StreamSpot 98.10% 99.00% 99.13% 99.27% 99.20% 0.90% 98.24% 99.77% 99.47% 99.26% 99.36% 1.53%
Unicorn Wget 97.34% 99.34% 99.23% 99.15% 99.19% 2.00% 98.00% 99.36% 99.66% 99.50% 99.58% 1.36%
DARPA CADETS 97.20% 99.03% 94.87% 99.78% 97.26% 1.83% 97.03% 99.15% 94.13% 99.71% 96.84% 2.12%
DARPA TRACE 96.03% 97.45% 95.22% 99.34% 97.23% 1.42% 96.18% 97.60% 93.28% 99.35% 96.24% 1.42%
DARPA THEIA 96.93% 98.94% 95.94% 99.40% 97.65% 2.01% 97.29% 99.23% 94.70% 99.22% 96.94% 1.94%
4.2. Overall Detection Efficacy Comparsion(Q1)
Table 2 and Table 3 show Winemaking ’s performance across various datasets using different teacher models. We observe that Winemaking ’s student model consistently improves ACC after distillation from all teacher models. Table 4 compares Winemaking with the baselines across multiple datasets. For StreamSpot, we compare Winemaking with StreamSpot, ThreaTrace, and MAGIC; for Unicorn Wget, we compare Winemaking with Unicorn, Prov-Gem, ThreaTrace, and MAGIC; for DARPA E3 datasets, we compare Winemaking with Prov-Gem, ThreaTrace, Log2vec, DeepLog, and MAGIC. Results indicate that Winemaking achieves the best or second-best results on all datasets. Specifically, on the StreamSpot, which involves relatively simple attack scenario graphs, Winemaking —being an entity-level detector—handles it quite easily. In the Unicorn Wget, which is somewhat more complex compared to StreamSpot but still graph-level, Winemaking only slightly lags behind Prov-Gem in Precision (PR), but outperforms all other baselines across all other metrics. In the three DARPA E3, Winemaking ranks second only to MAGIC in the Theia and Trace datasets, slightly behind in PR and F1-score.
These results demonstrate that Winemaking ’s knowledge distillation approach effectively transfers knowledge from complex GNN teacher models to the student model, enabling the student model to retain high anomaly detection accuracy across multiple detection mechanisms.
Table 4. Winemaking ’ comparison results.
Datasets Systems ACC PR RC F1
StreamSpot 66% 73% 91% 81%
ThreaTrace 96% 95% 93% 96%
MAGIC 99% 99% 99% 99%
StreamSpot Winemaking 99% 99% 100% 99%
Unicorn 90% 86% 95% 90%
Prov-Gem NA 100% 80% 89%
ThreaTrace 98% 93% 98% 95%
MAGIC 99% 98% 96% 97%
Unicorn Wget Winemaking 99% 99% 99% 99%
Log2vec 98% 49% 85% 62%
DeepLog 95% 23% 74% 35%
ThreaTrace 99% 90% 99% 95%
MAGIC 99% 94% 99% 97%
DARPA CADETS Winemaking 99% 95% 99% 97%
Log2vec 97% 54% 78% 64%
DeepLog 96% 41% 68% 51%
ThreaTrace 98% 72% 99% 83%
MAGIC 99% 99% 99% 99%
DARPA TRACE Winemaking 98% 95% 99% 97%
Log2vec 99% 62% 66% 64%
DeepLog 98% 16% 14% 15%
ThreaTrace 99% 87% 99% 93%
MAGIC 99% 98% 99% 98%
DARPA THEIA Winemaking 100% 96% 99% 98%
  • -   Best performance
  • -   Second best performance
4.3. Ablation Study (RQ2)
We conducted an ablation study on the CADETS dataset using GCN as the teacher model to investigate the individual contributions of Winemaking ’s different submodules, as shown in Figure 4.
[…]
((a)) FT
((b)) PRL
Figure 4. Ablation study
((a)) Embedding Dimension
((b)) Training Ratio
Figure 5. Hyperparameter Investigation
4.4. Hyperparameter Investigation (RQ3)
We conducted the following experiments to explore how key hyperparameters affect ACC. Using GCN as the teacher model, we tested the entire dataset.
Embedding Dimension: The embedding dimension determines the length of the vector representing each node. We tested dimensions 8, 16, 32, 64, 128. Figure 5.(a) shows that lower dimensions result in poorer performance due to limited expressiveness. The optimal performance was achieved at 32 dimensions. Higher dimensions did not improve performance and even slightly decreased it due to redundant features, increasing overfitting risk.
Training Ratio: The training ratio determines the number of labeled nodes in each batch. To further demonstrate the effectiveness of Vodka, we tested values of 5, 10, 20, 50, 100. The experimental results are shown in Figure 5.(b). As we can observe, the classification performance of the student model improves as the number of labeled nodes increases, reaching near-optimal performance when the batch contains 100 labeled nodes. We attempt to analyze the reasons behind this phenomenon: Increasing the number of labeled nodes introduces more supervised signals, which in turn enhances the efficiency of label inference during PRL. For a semi-supervised system like Winemaking, this improvement in label propagation is undoubtedly significant in boosting the model’s performance.
4.5. Performance Overhead (RQ4)
Winemaking aims to perform APTs detection with minimal overhead and flexible deployment in various environments. In real-world deployments, GPUs may be unavailable, and only CPUs may be used. Therefore, we primarily focused on Winemaking ’s detection time and memory usage under CPU-only conditions. Table 5 compares Winemaking ’s performance overhead with ShadeWatcher and MAGIC on the DARPA E3 TRACE dataset. Results show that Winemaking has significant advantages in both detection time and memory usage. At the same training ratio, Winemaking is 1.4 times faster than ShadeWatcher and 5.2 times faster than MAGIC.
Given its high detection speed, Winemaking could be deployed as a real-time online detector. For example, with the TRACE dataset, which took two weeks to collect, generating 1.37GB of audit logs per day, Winemaking only requires 23 seconds per day to complete the APTs detection. This means Winemaking ’s lightweight and high-speed detection capabilities make it suitable for daily APTs monitoring in small organizations and enterprises.
Figure 6. Robustness against adersarial attacks study.
[…]
A line figure showing the performance change with respect to the number of centric nodes, the masking subgraph scale, and the self-contrastive weight.
4.6. Robustness Against Adversarial Attacks (RQ5)
As more works focus on provenance-based data analysis, adversarial attacks(Chakraborty et al., 2021; Tramer and Boneh, 2019) designed to evade such provenance-based detectors are also increasing. One common adversarial attack is adversarial mimicry, which alters provenance data and adds benign representations to mimic benign behavior, disguising malicious system entities to evade detection. In this section, we simulate mimicry attacks to evaluate Winemaking ’s robustness against adversarial mimicry, using evaluation standards from (Goyal et al., 2023) and (Rehman et al., 2024). Our mimicry attack steps are as follows: first, we select benign system entities from the normal training data. Next, we introduce events and combine them with the benign entities. Finally, we embed the structure of the benign nodes into the attack graph. We compare Winemaking (using GCN as the teacher model and CADETS as the dataset) with FLASH(Rehman et al., 2024). Results (Figure 6) show that, as false events increase, FLASH’s anomaly score significantly decreases, indicating its weak robustness against mimicry attacks. In contrast, Winemaking maintains an anomaly score above 95% for all levels off false events, demonstrating its superior robustness.
The reasons for this are twofold: first, Winemaking is based on entity-level detection, making it highly sensitive to large-scale false event additions. Second, the provenance graph’s neighbor denoising process smooths out some of the strong malicious expressions, making it less susceptible to manipulation.
Table 5. Winemaking ’ performance overhead.
System Granularity
Time
consumpution(s)
Peak Memory
consumption (MB)
ShadeWatcher Edge 220 NA
MAGIC Node 825 1,667
Winemaking Node 158 982
5. Related Work
APTs Detection. Current research primarily focuses on provenance-based methods to detect APTs, which can be categorized into three types based on the way audit logs are utilized: statistical-based methods, rule-based methods, and learning-based methods. Statistical-based methods, such as Nodoze(Hassan et al., 2019), Swift(Hassan et al., 2020b), and PRIOTRACKER(Liu et al., 2018), quantify the rarity of statistical elements within the provenance graph, assigning suspiciousness scores to these elements, ultimately leading to anomaly detection. Rule-based methods, like Holmes(Milajerdi et al., 2019), SLEUTH(Hossain et al., 2017), and (Hassan et al., 2020a), collect prior attacks to help security personnel build attack rule libraries, enabling the detection of unknown attacks by matching them to predefined rules. Learning-based methods utilize machine learning techniques to model benign system behavior or attack patterns in order to identify unknown anomalous behaviors. Relevant work in this category includes (Jia et al., 2024; Zengy et al., 2022; Cheng et al., 2023; Wang et al., 2022; Han et al., 2020; Rehman et al., 2024).
Knowledge Distillation. Knowledge distillation was first proposed by (Hinton et al., 2015) and aims to transfer knowledge from a complex model (referred to as the teacher model) to a smaller model (referred to as the student model). This process enables the student model to reduce time and space complexity without compromising prediction quality, thus simplifying model inference and enhancing performance. While knowledge distillation has been widely applied in the field of computer vision, recent studies have extended the basic framework to other domains. Some works have also focused on the design of the student model. For example, (Yang et al., 2021) was the first to design a student model that is non-GCN-based, while (Tian et al., 2019) introduced contrastive learning into the distillation process to improve the student model’s ability to learn representations.
6. Conclusion
In this paper, we innovatively propose a novel APTs detection method called Winemaking, which integrates knowledge distillation and provenance-based APTs detection. Winemaking introduces a graph Laplacian-based approach for neighbor denoising and signal smoothing on provenance graphs. Additionally, Winemaking designs a knowledge distillation framework that distills GNNs knowledge from a complex teacher model into a lightweight student model. The student model combines feature transformation and label propagation, allowing for low-cost detection after training and reconstruction of the attack chain. We evaluate Winemaking’s performance on three widely used datasets, demonstrating excellent detection results.
Appendix A Appendix
A.1. Provenance Graph Description
Table 6 provides a description of the system event types and the edge relationships in the provenance graph.
System event type Relation Description
Process $\rightarrow$ R1 $\rightarrow$ Process ”R1”: ”fork”, ”execute”, ”exit”, ”clone”, ”change”,etc.
Process $\rightarrow$ R2 $\rightarrow$ File ”R2”: ”read”, ”open”, ”close”, ”write”, ”rename”,etc.
Process $\rightarrow$ R3 $\rightarrow$ Netflow ”R3”: ”connect”, ”send”, ”recv”, ”read”, ”close”,etc.
Process $\rightarrow$ R4 $\rightarrow$ Memory ”R4”: ”read”, ”mprotect”, ”mmap”, etc.
Table 6. System event and Rlation description.
A.2. Datasets
We provide a detailed description of the three datasets used(Table 7, Table 8, Table 9), including their specific contents and original sizes.
Scenarios # of Graphs
Average # of
Nodes
Average # of
Edges
YouTube 100 8,292 113,229
Gmail 100 6,826 37,382
Video Game 100 8,636 112,958
Download 100 8,830 310,814
CNN 100 8,989 294,903
Attack 100 8,890 28,423
Table 7. Overview of StreamSpot Dataset.
Scenarios
# Log
pieces
Average # of
Entity
Avg # of
Interaction
# of Raw
DataSize(GB)
Benign 125 265,424 975,226 64.0
Attack 25 257,156 957,968 12.6
Table 8. Overview of Unicorn Wget Datasets.
Datasets # of Node #of Edges
# of Raw Data
Size(GiB)
E3-THEIA 1,623,966 2,874,821 17.9
E3-CADETS 1,627,035 3,303,264 18.3
E3-TRACE 3,288,676 4,080,457 15.4
Table 9. Overview of Darpa-E3 Datasets.
A.3. Initial Parameters for Teacher and Student Models
We provide more information about the initial settings for both the teacher and student models.
Teacher Models are as follows:
  • GCN: 2 layers, 64 hidden units, learning rate of 0.01, dropout probability of 0.8, learning rate decay of 0.01.
  • GAT: 2 layers, 8 attention heads, 64 hidden units, learning rate of 0.01, dropout probability of 0.6, learning rate decay of 0.01, attention dropout probability of 0.3.
  • SAGE: 128 hidden units, learning rate of 0.01, sample size of 5, batch size of 256, learning rate decay of 0.005.
  • SGC: 2 layers, learning rate of 0.01, learning rate decay of 0.01.
Student Model: The number of layers $K$ is set to 5, the MLP hidden layer has 16 units, dropout rate is 0.2, the Adam optimizer has a learning rate of 0.01, and weight decay is set to 0.001.
A.4. Interpretability Analysis of the Student Model
We conduct an interpretability analysis of the student model. Recall that Vodka’s student model consists of two main mechanisms: the Feature Transformation (FT) mechanism, powered by MLP, and the Personalized PageRank Random Walk Label Propagation (PRL) mechanism. The label propagation mechanism includes a ”jump-back” feature, meaning there is a probability of returning to the previous node, which is controlled by the parameter $\alpha$. Additionally, the final prediction result of the student model is $f_{std}^{k+1}(v)=\beta_{v}f_{FT}(v)+(1-\beta_{v})f_{\mathrm{PRL}}^{k+1}(v)$, and this is controlled by the balancing parameter $\beta$. Our goal is to explore how these two parameters affect the bias of the student model’s prediction results. The experiments were conducted with a GCN as the teacher model and using the DARPA CADETS dataset.
Callback Probability $\alpha$: We dynamically adjust $\alpha$ and plot the neighborhood of the target node when $\alpha$ is at its maximum(Node 1) and minimum(Node 2) values, using different colors to represent labels. As shown in Figure 7, the neighbors of Node 1 are more diverse and numerous compared to those of Node 2. This indicates that nodes with a high $\alpha$ have a higher probability of jumping back to themselves, limiting the diversity of learning from neighboring nodes.
Balancing Parameter $\beta$: Similarly, we dynamically adjust $\beta$ and plot the neighborhood of the target node when $\beta$ is at its maximum(Node 3) and minimum(Node 4) values, with different colors representing labels. As shown in Figure 8, the labels of Node 4’s neighborhood are almost identical, whereas those of Node 3 are almost entirely different. This reflects that the Personalized PageRank Random Walk (PRL) mechanism contributes more to node prediction compared to FT, as when the balancing parameter favors PRL, the prior knowledge from labels and structure plays a more significant role in improving prediction accuracy.
((a)) Node 1
((b)) Node 2
Figure 7. Study of $\alpha$
((a)) Node 3
((b)) Node 4
Figure 8. Study of $\beta$